OpenStudy (bojack97):

I need help. If (5^4)^m = 5^12, what is the value of m?

hartnn (hartnn):

can you factor 12?? 12 = 4* ???

hartnn (hartnn):

\((A^b)^c = A^{bc}\) then use this property

OpenStudy (bojack97):

I don't under stand.

OpenStudy (bojack97):

wait, I add b and c?

OpenStudy (bojack97):

can you tell me the answer then explain how you got there?

OpenStudy (himansh):

ya man its so easy 5^4^m= 5^4m=5^12 so 4m =12 m =3

OpenStudy (bojack97):

oh ok thank you!

hartnn (hartnn):

from that property we get \((5^4)^m = 5^{4m} = 5^{12}\) so 4m = 12 get m from here

OpenStudy (himansh):

welcm :)